a body stands on a scale in the elevator if the scale reads 627.2 when the elevator is stationary, what will it read when the elevator is moving upward with an acceleration of 3.0m/s, a constant speed and a deceleration of 2m/s
1
Expert's answer
2020-12-14T07:30:50-0500

Putting "m=627.2kg"

scale reading when the elevator is moving upward with an acceleration of "3ms^{-2} = m(a+g)"

"=627.2(3+9.8)=627.2\times 12.8=8028.16N"

At constant speed, "a=0ms^{-2}"

"reading =m(a+g)=627.2(0+9.8) =6146.56N"

At a deceleration of "2ms^{-2}," "reading=627.2(-2+9.8)=627.2\times 7.8=4892.16N"